On November 15, 2015, Shields Company purchased inventory costing $12,400 on credit. The credit terms were 1/10, n//30.

a. Assume that Shields Company paid the invoice on November 23, 2015. Prepare journal entries to record the purchase of this inventory and the cash payment to the supplier using the net-of-discount method.

General Journal
Date Description Debit Credit
11/15 Inventory Answer

12,276

Answer

0

Accounts Payable Answer

0

Answer

12,276

11/23 Accounts Payable Answer

12,276

Answer

0

Cash Answer

0

Answer

12,276


b. Set up the necessary T-accounts and post the journal entries from question a. to the accounts.

Cash
Answer

0

Answer

12,276

Inventory
Answer

12,276

Answer

0

Accounts Payable
Answer

12,276

Answer

12,276


c. Compute the cost of a lost discount as an annual percentage rate. Round answer to one decimal place (i.e., 34.6%).

Answer

%
Answer #1

c. Cost of a lost discount = [D/(1 – D)] x [365/(F – L)]

Where, D = 0.01; F = 30 days; L = 10 days

Cost of a lost discount = [0.01/(1 – 0.01)] x [365/(30 - 10)] = [0.01/0.99] x [365/20] = [0.01/0.99] x 18.25 = 18.4%